AddieroniAndCheese:

High key stressing over here SOMEONE HELP PLEASE DUE TONIGHT a) Write 2-3 sentences to describe whether it takes more energy to raise the temperature by one degree of a 1-g piece of gold or a 1-g piece of magnesium. Explain your reasoning. b) Identify the heat that is lost when 5.6 g of aluminum at 65.2 °C is cooled to 24.7 °C. c) If 5.6 g of aluminum at 65.2 °C is placed in a calorimeter that contains water at 23.5 °C and the temperature of the water after the heat exchange occurs is 24.7 °C, then identify how much water is in the calorimeter.

AddieroniAndCheese:

@gdubb08 :D

AddieroniAndCheese:

1 attachment
gdubb08:

It takes more energy to raise the temperature of a 1-g piece of magnesium by one degree compared to a 1-g piece of gold. This is because magnesium has a higher specific heat capacity (1024 J/(kg·K)) than gold (130 J/(kg·K))[1][3]. The higher specific heat capacity of magnesium means it requires more energy to increase its temperature by the same amount as gold. b) The heat lost by the aluminum can be calculated using the formula q = m × c × ΔT. For aluminum, c = 887 J/(kg·K)[1], m = 5.6 g = 0.0056 kg, and ΔT = 24.7°C - 65.2°C = -40.5°C. Therefore, q = 0.0056 × 887 × (-40.5) = -200.7 J. The negative sign indicates heat is lost from the aluminum. c) To find the mass of water in the calorimeter, we can use the principle that heat lost by aluminum equals heat gained by water. Using q = m × c × ΔT for water, where c = 4184 J/(kg·K)[5], ΔT = 24.7°C - 23.5°C = 1.2°C, and q = 200.7 J (heat gained by water), we can solve for m: 200.7 = m × 4184 × 1.2. This gives us m = 0.0400 kg or 40.0 g of water in the calorimeter.

AddieroniAndCheese:

THANK YOU!!!!

gdubb08:

@addieroniandcheese wrote:
THANK YOU!!!!
np

gdubb08:

@addieroniandcheese wrote:
THANK YOU!!!!
any more?

AddieroniAndCheese:

no, I don't think so, yet, I have one more assignment but I don't know how that's gonna be

gdubb08:

@addieroniandcheese wrote:
no, I don't think so, yet, I have one more assignment but I don't know how that's gonna be
alright. Lemme know if u need help at all with anything. Ima head 2 bed in a few or sum. I'll be bk 2morrow maybe, or 2sday.

AddieroniAndCheese:

kk thank you!!

AddieroniAndCheese:

its end of marking period and I idiocally procrastinated lol